OpenStudy (anonymous):

A sled is pushed along an ice covered lake. It has some initial velocity before coming to a rest in 15 m. It took 23 seconds before the sled and rider come to a rest. If the rider and sled have a combined mass of 52.5 kg, what is the magnitude and direction of the stopping force? What do "we" call the stopping force?

OpenStudy (anonymous):

You are trying to find the deceleration (-a) so you can use: F = ma to find the 'stopping' force. From these equations: V2 = V1 + at X2 = X1 + V1t + (1/2) at^2 you can get an equation for V1 from the first equation, put it into the second equation to find the acceleration.

OpenStudy (anonymous):

if i use the equation V2=V1+at, how will i find V1 if a is also unknown?

OpenStudy (anonymous):

V2 = 0 V1 = -at You can use that for V1 in the second equation and solve for a.

OpenStudy (anonymous):

oh, I got it. Thanks!